OpenStudy (anonymous):

Ventura Coal mined 6 ⅔ tons on Monday, 7 ¾ tons on Tuesday, and 4 ½ tons on Wednesday. If the goal is to mine 25 tons this week, how many more tons must be mined?

OpenStudy (dumbcow):

i believe this is just simple addition add up the mixed numbers....then subtract from 25 6+7+4 = 17 \[\frac{2}{3}+\frac{3}{4}+\frac{1}{2} = \frac{8}{12}+\frac{9}{12}+\frac{6}{12} = \frac{23}{12}\] total: 18 11/12 25 - 18 11/12 = 6 1/12

OpenStudy (anonymous):

thanks